OpenStudy (anonymous):

FAN AND MEDAL!! How many grams of potassium chloride must decompose to produce 69.9 milliliters of potassium metal, if the density of potassium is 1.78 g/mL? Show all steps of your calculation as well as the final answer. KCl → K + Cl2

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

1/ 2KCl → 2K + Cl2 one molecule of KCl will give one molecule of K so their amounts are equal n(KCl) = n(K) , n(K) = m(K)/M(K), m(K) = r(density) * V so n(K) = (r(density) * V)/M(K) m(KCl) = n(KCl) * M(KCl) m(KCl) = [(r(density) * V)/M(K)] * M(KCl) m(KCl) = [ (69.9 mL * 1.78g/mL)/39 g/ mol] * 74.55 g mol−1 m(KCl) = 237.83 g

OpenStudy (anonymous):

hopefully that helps

OpenStudy (anonymous):

Wow! Thank you so much man! The final answer is m(KCl) = 237.83 g??

OpenStudy (anonymous):

yes it is and your welcome. Plus im a girl

OpenStudy (anonymous):

Lol I call everyone man

OpenStudy (anonymous):

@nacholover

OpenStudy (anonymous):

oh ok.But your welcome anything else i can help you with.

OpenStudy (anonymous):

Yes gimme a few moments I have 2 more chem questions!! @nacholover

OpenStudy (anonymous):

k just tag me in the post so i can answer them